# Scalable Estimation of Multinomial Response Models

### v4

@@ 6.1 Empirical Results
Consideration sets No No $K=\infty$ $K=\infty$ $K=1$ $K=1$
@@ 7 Discussion
In this concluding section, we discuss the broader relevance of the work, especially to the modeling of excess zeros in high-dimensional sparse microbiome data sets. In a microbiome dataset with $n$ samples and $J$ taxa, let $x_{ij}$ denote the measured count for taxon $j$ in sample $i$, and $T_{i}=\sum_{j=1}^{J}x_{ij}$ represent the total count for all taxa in the $i$th sample, where $i=1,\ldots,n$ and $j=1,\ldots,J$. Typically it is assumed that the vector $\bm{x}_{i}=(x_{i1},\ldots,x_{iJ})^{\prime}$ follows a multinomial distribution with index $T_{i}$ and a vector of probabilities $\bm{\rho}_{i}=(\rho_{i1},\ldots,\rho_{iJ})^{\prime}$, where $10<\rho_{ij}<$ 1 and $\sum_{j=1}^{J}\rho_{ij}=1$. In our notation, $Y_{it}$ denotes the selected item and relates to $\bm{x}_{i}$ through $x_{ij}=\sum_{t=1}^{T_{i}}1(Y_{it}=j)$. To address the high dimensionality and sparsity of such datasets, Zeng et al. (2023) propose the following hierarchical model translated in our terminology as
[…]
where $1-C_{ij}$ are latent indicators for excess zeros, and $q_{j}$ are the corresponding probabilities. The $\bm{f}_{i}$ are latent factors, and the $\bm{\beta}_{j}$ denote the loadings of the associated factors. Note that the excess zeros across taxa are independent in this modeling. In other words, their model corresponds to the independent consideration model that we review in the introduction. Our modeling can be used if excess zeros exhibit complex dependency patterns across taxa. In this more general setting, $\bm{C}_{i}=(C_{i1},\ldots,C_{iJ})^{\prime}$ are correlated vectors and i.i.d. with the density for the infinite mixture of independent consideration models (3). Specifically, we would have $\Pr(\bm{C}_{i}=\bm{c}_{i})=\sum_{h=1}^{\infty}\omega_{h}\prod_{j=1}^{J}\left\{q_{hj}^{c_{ij}}\left(1-q_{hj}\right)^{1-c_{ij}}\right\}$. We can adapt the MCMC framework of this paper to estimate this model.
Another key issue in practice is variable selection when many subject-level covariates are available. This challenge can be addressed using shrinkage priors. A natural extension of our framework involves modeling consideration sets that change at one or two points in time due to learning from past choices. This would require incorporating the learning process into the model, and modifying the theoretical analysis accordingly. We leave this promising direction for future work. Matlab code for estimating our model is available.
@@ Appendix A Proof of Lemma 2
Acknowledgment
We benefited from useful comments from Dimitris Korobilis, Paola Manzini, Francesca Molinari, Nail Kashaev, Alessandro Iaria, Ao Wang, Elena Manresa, Martin Weidner, Laura Liu, Asim Ansari, Andriy Norets, Toru Kitagawa, Frank Schorfheide, Zhentong Lu, Victor Aguiar, Yao Luo, Victor Aguirregabiria, Andrew Ching, Kosuke Uetake, Matthijs Wildenbeest, Mitsuru Igami, and the participants at the economics and statistics seminars at the University of Alberta, the 2023 NBER-NSF Seminar in Bayesian Econometrics and Statistics (SBIES) at the Federal Reserve Bank of Philadelphia, the 2023 INFORMS Annual Meeting in Phoenix, and the 2024 International Industrial Organization Conference in Boston. We gratefully acknowledge financial support from the Canadian Social Sciences and Humanities Research Council Insight Development Grant. Disclaimer: Researchers’ own analyses calculated (or derived) based in part on data from Nielsen Consumer LLC and marketing databases provided through the NielsenIQ Datasets at the Kilts Center for Marketing Data Center at The University of Chicago Booth School of Business. The conclusions drawn from the NielsenIQ data are those of the researchers and do not reflect the views of NielsenIQ. NielsenIQ is not responsible for, had no role in, and was not involved in analyzing and preparing the results reported herein.
@@ Acknowledgment
Section SB provides intermediate results used to prove Theorems 1 and 2 of the main paper, and their proofs. Section SC presents the conditional posterior distributions of the parameters other than the consideration sets. Section SD illustrates the impact of the prior choice for the attention probabilities on the prior on the distribution of consideration sets. Section SE shows additional simulation results under $J=4$. Section SF provides additional results from the empirical application.
@@
Suppose $\mathcal{W}$ is compact. Let $(K,\bm{\phi}_{1:K})$ and $\bm{w}\in\mathcal{W}$. Then for each $j\in\mathcal{J}$, $\forall\varepsilon>0$ and $\bm{\theta}^{(1)}=\{\bm{\beta}^{(1)},\bm{D}^{(1)}\}$, $\exists\delta>0$ such that for any $\bm{\theta}^{(2)}=\{\bm{\beta}^{(2)},\bm{D}^{(2)}\}$ satisfying $||\bm{\beta}^{(1)}-\bm{\beta}^{(2)}||<\delta$ and
@@
$\left|p(j|\bm{w};\bm{\theta}^{(1)},K,\bm{\phi}_{1:K})-p(j|\bm{w};\bm{\theta}^{(2)},K,\bm{\phi}_{1:K})\right|$ $\leq\sum_{c\in\mathcal{C}}\pi(c|K,\bm{\phi}_{1:K})\left|\Pr(j|\bm{\theta}^{(1)},\bm{w},c)-\Pr(j|\bm{\theta}^{(2)},\bm{w},c)\right|$
$=\sum_{c:j\in c}\pi(c|K,\bm{\phi}_{1:K})\left|\Pr(j|\bm{\theta}^{(1)},\bm{w},c)-\Pr(j|\bm{\theta}^{(2)},\bm{w},c)\right|,$
@@
$|\displaystyle\left|\Pr(j|\bm{\theta}^{(1)},\bm{w},c))-\Pr(j|\bm{\theta}^{(2)},\bm{w},c)\right|$$\leq\left|\Pr(j|\{\bm{\beta}^{(1)},\bm{D}^{(1)}\},\bm{w},c)-\Pr(j|\{\bm{\beta}^{(2)},\bm{D}^{(1)}\},\bm{w},c)\right|$ (SB.1)
$+\left|\Pr(j|\{\bm{\beta}^{(2)},\bm{D}^{(1)}\},\bm{w},c)-\Pr(j|\{\bm{\beta}^{(2)},\bm{D}^{(2)}\},\bm{w},c)\right|.$ (SB.2)
@@
$\leq\sum_{c}\pi_{c}\left|\prod_{t=1}^{T}\Pr\left(Y_{it}=y_{t}|\bm{\beta},\bm{w}_{t},c)-\prod_{t=1}^{T}\Pr\left(Y_{it}=y_{t}|\bm{\beta}^{\prime},\bm{w}_{t},c)\right|$
@@
$\leq\gamma_{1}||\bm{\beta}-\bm{\beta}^{\prime}||_{2}+\gamma_{2}||\bm{\pi}-\bm{\pi}^{\prime}||_{1},$
@@
[…]
@@
Panel (a) in Figure SD.1 shows the implied prior distribution over the consideration sets under the uniform prior on $q_{hj}$ when $J=4$. Under the uniform prior, the prior expectation of $q_{hj}=0.5$, so the prior is the same across all consideration sets and is centered around $0.06250.5^{4}=$ 0.0625. Panel (b) gives results under our sparsity-supporting prior. In this case, the prior distribution shrinks to 00 as the cardinality of the consideration set increases.
@@
Figure SE.1: Histograms of estimated posterior probabilities of $H_{1}$ in each of the 100 simulations under (a) case 1 (dependent consideration - $H_{1}$ is true) and (b) case 2 (independent consideration - $H_{0}$ is true). $\varepsilon=0.1,n=50$. $T=5$.
We use the current data-generating process as the first case (dependent consideration). In the second case, the consideration is independent. We generate $C_{ij}\overset{iid}{\sim}\text{Bernoulli}(\gamma_{j})$, for $j=1,2,3$ with $(\gamma_{1},\gamma_{2},\gamma_{3})=(0.2,0.15,0.35)$ and fix $C_{i4}=1$, for $i=1,\ldots,n$. Figure SE.1 (a) provides a histogram showing the estimated posterior probabilities of $H_{1}$ under the first case ($H_{1}$ is true) across the 100 simulated data sets using $\varepsilon=0.1$. The method appropriately assigns a value close to one to $\Pr(H_{1}|\bm{D}^{n})$ in most cases, with only $3/100$ having an estimated $\Pr(H_{1}|\bm{D}^{n})<0.5$. Figure SE.1 (b) provides the results for case 2. The posterior probability assigned to $H_{1}$ is close to zero for most simulations. We find similar results with random effects, as shown below.
@@
We repeat the simulation study now with preference heterogeneity. We generate the consideration sets as before, but now use the random effects logit with the specification $V_{ijt}=\delta^{*}_{j}+(\beta^{*}+b_{i})x_{ijt},$ where $b_{i}\sim N(0,D^{*})$ with $D^{*}=1$. We fit the random effects logit with the proposed flexible approach for the distribution of consideration sets. The results are presented in Table SE.1. We see that as $n$ increases, the MSEs/bias/standard deviation tend to decrease. However, this is not the case when the independent consideration is imposed, i.e. $K=1$. Also, MSEs/biases are larger in general than for the proposed flexible approach. In addition, there are distortions of the coverage of the credible intervals under $K=1$. Figure SE.2 shows that for $K=\infty$, we see the posterior on $\bm{\pi}$ approaches to the truth while it does not under $K=1$, due to the mis-specification.
Table SE.1: Simulation results for $J=4$ with random effects
$(K,T)$ $n$ $\beta$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $D^{1/2}$ $\pi$ Time
$(\infty,5)$ $50$ 0.135 ( 0.136 ; 0.44 ) 0.229 ( 0.418 ; 0.92 ) 0.249 ( 0.421 ; 0.92 ) 0.149 ( 0.321 ; 0.92 ) 0.722 ( 0.064 ; 0.06 ) 0.444 ( 0.03 ) 2.36
$500$ 0.009 ( 0.079 ; 0.90 ) 0.019 ( 0.151 ; 0.96 ) 0.021 ( 0.149 ; 0.96 ) 0.014 ( 0.123 ; 0.90 ) 0.017 ( 0.093 ; 0.84 ) 0.211 ( 0.01 ) 24.86
$(\infty,15)$ $50$ 0.037 ( 0.159 ; 0.86 ) 0.046 ( 0.211 ; 0.96 ) 0.039 ( 0.207 ; 0.96 ) 0.026 ( 0.182 ; 0.96 ) 0.091 ( 0.154 ; 0.72 ) 0.358 ( 0.01 ) 3.11
$500$ 0.003 ( 0.059 ; 0.96 ) 0.005 ( 0.067 ; 0.96 ) 0.004 ( 0.067 ; 0.96 ) 0.004 ( 0.058 ; 0.98 ) 0.004 ( 0.055 ; 0.92 ) 0.129 ( 0 ) 37.7
$(1,5)$ $50$ 0.156 ( 0.13 ; 0.30 ) 0.732 ( 0.365 ; 0.48 ) 0.925 ( 0.369 ; 0.44 ) 0.142 ( 0.318 ; 0.86 ) 0.733 ( 0.058 ; 0.01 ) 0.744 ( 0.03 ) 2.14
$500$ 0.032 ( 0.068 ; 0.36 ) 1.027 ( 0.129 ; 0.01 ) 1.145 ( 0.13 ; 0.01 ) 0.016 ( 0.119 ; 0.90 ) 0.035 ( 0.082 ; 0.50 ) 0.719 ( 0.01 ) 21.85
$(1,15)$ $50$ 0.037 ( 0.159 ; 0.84 ) 0.082 ( 0.223 ; 0.86 ) 0.074 ( 0.22 ; 0.88 ) 0.032 ( 0.186 ; 0.96 ) 0.09 ( 0.149 ; 0.68 ) 0.393 ( 0.01 ) 2.86
$500$ 0.003 ( 0.058 ; 0.96 ) 0.022 ( 0.072 ; 0.60 ) 0.02 ( 0.071 ; 0.62 ) 0.007 ( 0.06 ; 0.82 ) 0.003 ( 0.054 ; 0.92 ) 0.207 ( 0.01 ) 33.89
  • For $\beta$, $\delta$, and $D^{1/2}$, we show the average of the mean squared errors (MSEs), using the posterior means as point estimators. In parentheses, the average posterior standard deviations (SDs) and empirical coverages of 95% credible intervals are shown. For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ with the posterior SDs in the parentheses. Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM.
@@
Figure SE.3: Histograms of estimated posterior probabilities of $H_{1}$ in each of the 100 simulations under (a) case 1 (dependent consideration - $H_{1}$ is true) and (b) case 2 (independent consideration - $H_{0}$ is true). $\varepsilon=0.1,n=50$. $T=5$.

### v5

@@ 6.1 Empirical Results
Consideration sets No No $\infty K=K=\infty$ $K=1$ $K=1$
@@ 7 Discussion
In this concluding section, we discuss the broader relevance of the work, especially to the modeling of excess zeros in high-dimensional sparse microbiome data sets. In a microbiome dataset with $n$ samples and $J$ taxa, let $u_{ij}$ denote the measured count for taxon $j$ in sample $i$, and $T_{i}=\sum_{j=1}^{J}u_{ij}$ represent the total count over taxa in the $i$th sample, where $i=1,\ldots,n$ and $j=1,\ldots,J$. Typically it is assumed that the vector $\bm{u}_{i}=(u_{i1},\ldots,u_{iJ})^{\prime}$ follows a multinomial distribution with index $T_{i}$ and a vector of probabilities $\bm{\rho}_{i}=(\rho_{i1},\ldots,\rho_{iJ})^{\prime}$, where $0<\rho_{ij}<1$ and $\sum_{j=1}^{J}\rho_{ij}=1$. In our notation, $u_{ij}$ relates to $Y_{it}$ through $u_{ij}=\sum_{t=1}^{T_{i}}1(Y_{it}=j)$. To address the high dimensionality and sparsity of such datasets, Zeng et al. (2023) propose the following hierarchical model translated in our terminology as
$\Pr(\bm{C}_{i}=\bm{c}_{i})=q_{j}^{c_{ij}}\left(1-q_{j}\right)^{1-c_{ij}},\quad f_{i1},\ldots,f_{ik}\overset{\text{ind}}{\sim}\mathcal{N}(0,1),$
$\bm{u}_{i}|\bm{\rho}_{i},T_{i}\overset{\text{ind}}{\sim}\text{MN}(\bm{\rho}_{i},T_{i}),\quad\rho_{ij}=\frac{C_{ij}\exp\left(\beta_{0j}+\bm{f}_{i}^{\prime}\bm{\beta}_{j}\right)}{\sum_{\ell=1}^{J}C_{i\ell}\exp\left(\beta_{0\ell}+\bm{f}_{i}^{\prime}\bm{\beta}_{\ell}\right)},$
where $\bm{C}_{i}=(C_{i1},\ldots,C_{iJ})^{\prime}$, and $1-C_{ij}$ are latent indicators for excess zeros, and the $q_{j}$ are the corresponding probabilities. The $\bm{f}_{i}$ are latent factors, and the $\bm{\beta}_{j}$ denote the loadings of the associated factors. Note that the excess zeros across taxa are independent in this modeling. In other words, the model above corresponds to the independent consideration model that we review in the introduction. In this context, complex dependency patterns across taxa in the excess zeros can be captured by our modeling. To do this, we would let $\bm{C}_{i}$ be correlated vectors and i.i.d. with the density for the infinite mixture of independent consideration models (3):
$\Pr(\bm{C}_{i}=\bm{c}_{i})=\sum_{h=1}^{\infty}\omega_{h}\prod_{j=1}^{J}\left\{q_{hj}^{c_{ij}}\left(1-q_{hj}\right)^{1-c_{ij}}\right\}.$
We can adapt the MCMC framework of this paper to estimate this model. Our approach for updating the $\bm{C}_{i}$’s and the mixture parameters can be used in conjunction with existing approaches for simulating the factor-related objects.
Another key issue in practice is variable selection when many subject-level covariates are available. This challenge can be addressed using shrinkage priors. A natural extension of our framework involves modeling consideration sets that change at one or two points in time due to learning from past choices. This would require incorporating the learning process into the model and modifying the theoretical analysis accordingly. We leave this promising direction for future work.
@@ Appendix A Proof of Lemma 2
Acknowledgments
We thank the editor and the review panel for excellent comments and suggestions that have improved the manuscript. We benefited from helpful comments from Dimitris Korobilis, Andriy Norets, Nail Kashaev, Alessandro Iaria, Ao Wang, Zhentong Lu, Toru Kitagawa, Yao Luo, Kosuke Uetake, Victor Aguiar, Laura Liu, Paola Manzini, Francesca Molinari, Elena Manresa, Martin Weidner, Asim Ansari, Frank Schorfheide, Victor Aguirregabiria, Andrew Ching, Matthijs Wildenbeest, Mitsuru Igami, and the participants at the economics and statistics seminars at the University of Alberta, the 2023 NBER-NSF Seminar in Bayesian Econometrics and Statistics (SBIES) at the Federal Reserve Bank of Philadelphia, the 2023 INFORMS Annual Meeting in Phoenix, the 2024 International Industrial Organization Conference in Boston, the 2025 Marketing Science Conference in Washington DC, and the 2025 Econometric Society World Congress in Seoul. Kenichi Shimizu gratefully acknowledges financial support from the Canadian Social Sciences and Humanities Research Council Insight Development Grant. Disclaimer: Researchers’ own analyses calculated (or derived) based in part on data from Nielsen Consumer LLC and marketing databases provided through the NielsenIQ Datasets at the Kilts Center for Marketing Data Center at The University of Chicago Booth School of Business. The conclusions drawn from the NielsenIQ data are those of the researchers and do not reflect the views of NielsenIQ. NielsenIQ is not responsible for, had no role in, and was not involved in analyzing and preparing the results reported herein.
@@ Acknowledgments
Section SB provides intermediate results used to prove Theorems 1 and 2 of the main paper, and their proofs. Section SC presents the conditional posterior distributions of the parameters other than the consideration sets. Section SD illustrates the impact of the prior choice for the attention probabilities on the prior on the distribution of consideration sets. Section SE shows additional simulation results. Section SF provides additional results from the empirical application.
@@
Suppose $\mathcal{W}$ is compact. Let $(K,\bm{\phi}_{1:K})$ and $\bm{w}\in\mathcal{W}$. Then for each $j\in\mathcal{J}$, $\forall\varepsilon>0$ and $\bm{\theta}^{(1)}=\{\bm{\beta}^{(1)},\bm{D}^{(1)}\}$, $\exists\delta>0$ such that for any $\bm{\theta}^{(2)}=\{\bm{\beta}^{(2)},\bm{D}^{(2)}\}$ satisfying $\bm{\beta}^{(1)}-\bm{\beta}^{(2)}||<\delta$ and
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\bm{\beta}^{(g+1)}$ $\text{ from }\bm{\beta}\big{|}\big{\{}y_{it}\big{\}},\big{\{}\bm{C}_{i}^{(g)}\big{\}},\bm{\delta}^{(g)},\big{\{}\bm{b}_{i}^{(g)}\big{\}},$
$\big{\{}\bm{b}_{i}^{(g+1)}\big{\}}$ $\text{ from }\big{\{}\bm{b}_{i}\big{\}}\big{|}\big{\{}y_{it}\big{\}},\big{\{}\bm{C}_{i}^{(g)}\big{\}},\bm{\delta}^{(g)},\bm{\beta}^{(g+1)},\bm{D}^{(g)},$
$\bm{D}^{(g+1)}$ $\text{ from }\bm{D}\big{|}\big{\{}\bm{b}_{i}^{(g+1)}\big{\}},$
$\bm{\delta}^{(g+1)}$ $\text{ from }\bm{\delta}\big{|}\big{\{}y_{it}\big{\}},\big{\{}\bm{C}_{i}^{(g)}\big{\}},\bm{\beta}^{(g+1)},\big{\{}\bm{b}_{i}^{(g+1)}\big{\}},$
$\big{\{}\bm{C}_{i}^{(g+1)}\big{\}}$ $\text{ from }\big{\{}\bm{C}_{i}\big{\}}\big{|}\big{\{}y_{it}\big{\}},\bm{\delta}^{(g+1)},\bm{\beta}^{(g+1)},\big{\{}\bm{b}_{i}^{(g+1)}\big{\}},\{\bm{q}_{h}^{(g)}\},\{S_{i}^{(g)}\},$
$\{V_{h}^{(g+1)}\}$ $\text{ from }\{V_{h}\}\big{|}\{S_{i}^{(g)}\},\alpha^{(g)},$
$\{\bm{q}_{h}^{(g+1)}\}$ $\text{ from }\{\bm{q}_{h}\}\big{|}\big{\{}\bm{C}_{i}^{(g+1)}\big{\}},\{S_{i}^{(g)}\},$
$\{u_{i}^{(g+1)}\}$ $\text{ from }\{u_{i}\}\big{|}\{S_{i}^{(g)}\},\{V_{h}^{(g+1)}\},$
$\{S_{i}^{(g+1)}\}$ $\text{ from }\{S_{i}\}\big{|}\{u_{i}^{(g+1)}\},\{\bm{q}_{h}^{(g+1)}\},\{V_{h}^{(g+1)}\},\big{\{}\bm{C}_{i}^{(g+1)}\big{\}},$
$\alpha^{(g+1)}$ $\text{ from }\alpha\big{|}\{V_{h}^{(g+1)}\},\{S_{i}^{(g+1)}\}.$
@@
Panel (a) in Figure SD.1 shows the implied prior distribution over the consideration sets under the uniform prior on $q_{hj}$ when $J=4$. Under the uniform prior, the prior expectation of $q_{hj}=0.5$, so the prior is the same across all consideration sets and is centered around $0.5^{4}=0.0625$. Panel (b) gives results under our sparsity-supporting prior. In this case, the prior distribution shrinks to 0 as the cardinality of the consideration set increases.
@@
Figure SE.1: Histograms of estimated posterior probabilities of $H_{1}$ in each of the 200 simulations under (a) case 1 (dependent consideration - $H_{1}$ is true) and (b) case 2 (independent consideration - $H_{0}$ is true). $\varepsilon=0.1,n=50$. $T=5$.
We use the current data-generating process as the first case (dependent consideration). In the second case, the consideration is independent. We generate $C_{ij}\overset{iid}{\sim}\text{Bernoulli}(\gamma_{j})$, for $j=1,2,3$ with $(\gamma_{1},\gamma_{2},\gamma_{3})=(0.2,0.15,0.35)$ and fix $C_{i4}=1$, for $i=1,\ldots,n$. Figure SE.1 (a) provides a histogram showing the estimated posterior probabilities of $H_{1}$ under the first case ($H_{1}$ is true) across the 200 simulated data sets using $\varepsilon=0.1$. The method appropriately assigns a value close to one to $\Pr(H_{1}|\bm{D}^{n})$ in most cases, with only $9/100$ having an estimated $\Pr(H_{1}|\bm{D}^{n})<0.5$. Figure SE.1 (b) provides the results for case 2. The posterior probability assigned to $H_{1}$ is close to zero for most simulations. We find similar results with random effects, as shown below.
@@
We repeat the simulation study now with preference heterogeneity. We generate the consideration sets as before, but now use the random effects logit with the specification $V_{ijt}=\delta^{*}_{j}+(\beta^{*}+b_{i})x_{ijt},$ where $b_{i}\sim N(0,D^{*})$ with $D^{*}=1$. We fit the random effects logit with the proposed flexible approach for the distribution of consideration sets. The results are presented in Table SE.1. We see that as $n$ increases, RMSEs/L1-errors/SDs tend to decrease. However, this is not the case when the independent consideration is imposed, i.e. $K=1$. Also, RMSEs/L1-errors are larger in general than for the proposed flexible approach. In addition, there are distortions of the coverage of the credible intervals under $K=1$. In Figure SE.2, for $K=\infty$, we see the posterior on $\bm{\pi}$ approaches to the truth while it does not under $K=1$, due to the mis-specification.
Table SE.1: Simulation results with $J=4$ (random effects)
$(K,T)$ $n$ $\beta$ $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ $D^{-1/2}$ $\pi$ Time
RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov RMSE (MCE)    SD (ESD)       Cov L1-error (MCE) SD (ESD)       Cov
$(\infty,5)$ $50$ 0.363 ( 0.011 ) 0.14 ( 0.195 ) 0.45 0.448 ( 0.025 ) 0.42 ( 0.445 ) 0.93 0.49 ( 0.028 ) 0.43 ( 0.462 ) 0.92 0.357 ( 0.018 ) 0.32 ( 0.289 ) 0.9 0.845 ( 0.006 ) 0.06 ( 0.133 ) 0.05 0.455 ( 0.009 ) 0.03 ( 0.028 ) 0.96 2.7
$300$ 0.124 ( 0.006 ) 0.1 ( 0.102 ) 0.85 0.221 ( 0.012 ) 0.19 ( 0.207 ) 0.93 0.212 ( 0.01 ) 0.19 ( 0.204 ) 0.95 0.151 ( 0.008 ) 0.16 ( 0.151 ) 0.97 0.215 ( 0.012 ) 0.12 ( 0.154 ) 0.72 0.268 ( 0.003 ) 0.02 ( 0.015 ) 0.93 20.97
$(\infty,15)$ $50$ 0.194 ( 0.01 ) 0.16 ( 0.183 ) 0.87 0.205 ( 0.01 ) 0.21 ( 0.205 ) 0.95 0.188 ( 0.01 ) 0.21 ( 0.187 ) 0.97 0.172 ( 0.009 ) 0.18 ( 0.17 ) 0.97 0.315 ( 0.014 ) 0.15 ( 0.198 ) 0.66 0.356 ( 0.005 ) 0.03 ( 0.022 ) 0.97 3.69
$300$ 0.075 ( 0.004 ) 0.07 ( 0.074 ) 0.95 0.086 ( 0.004 ) 0.09 ( 0.086 ) 0.95 0.079 ( 0.004 ) 0.09 ( 0.079 ) 0.96 0.066 ( 0.003 ) 0.07 ( 0.066 ) 0.99 0.081 ( 0.004 ) 0.07 ( 0.067 ) 0.9 0.193 ( 0.003 ) 0.01 ( 0.012 ) 0.91 57.4
$(1,5)$ $50$ 0.384 ( 0.011 ) 0.13 ( 0.184 ) 0.35 0.786 ( 0.024 ) 0.37 ( 0.463 ) 0.54 0.911 ( 0.029 ) 0.38 ( 0.491 ) 0.47 0.364 ( 0.018 ) 0.32 ( 0.303 ) 0.88 0.849 ( 0.005 ) 0.06 ( 0.1 ) 0.04 0.843 ( 0.006 ) 0.03 ( 0.025 ) 0.47 2.4
$300$ 0.192 ( 0.005 ) 0.09 ( 0.085 ) 0.47 1.00 ( 0.015 ) 0.16 ( 0.214 ) 0.01 1.058 ( 0.016 ) 0.17 ( 0.223 ) 0.00 0.166 ( 0.008 ) 0.15 ( 0.162 ) 0.94 0.261 ( 0.009 ) 0.1 ( 0.134 ) 0.48 0.862 ( 0.005 ) 0.01 ( 0.013 ) 0.00 19.7
$(1,15)$ $50$ 0.193 ( 0.01 ) 0.16 ( 0.181 ) 0.87 0.288 ( 0.016 ) 0.23 ( 0.246 ) 0.9 0.272 ( 0.018 ) 0.23 ( 0.23 ) 0.9 0.177 ( 0.01 ) 0.19 ( 0.177 ) 0.94 0.315 ( 0.014 ) 0.15 ( 0.2 ) 0.68 0.712 ( 0.002 ) 0.02 ( 0.018 ) 0.89 3.4
$300$ 0.076 ( 0.004 ) 0.07 ( 0.074 ) 0.96 0.159 ( 0.007 ) 0.09 ( 0.102 ) 0.72 0.15 ( 0.007 ) 0.09 ( 0.095 ) 0.77 0.085 ( 0.004 ) 0.08 ( 0.069 ) 0.92 0.078 ( 0.004 ) 0.07 ( 0.066 ) 0.93 0.704 ( 0.001 ) 0.01 ( 0.008 ) 0.63 56.98
  • For $\beta$, $\delta$, and $D^{1/2}$, for each case, we show the estimated root mean squared error (RMSE), using the posterior means as point estimator. In parenthesis, the jackknife estimate of Monte Carlo Error (MCE) for the RMSE is presented. Next, the average of the posterior standard deviations (SD) is shown with the empirical standard deviation (ESD) of the posterior mean in the parenthesis. Third, the empirical coverage (Cov) of 95% credible interval is given.
  • For $\pi$, we show the average of $L_{1}$ norm between the posterior mean and $\bm{\pi}^{*}$ (L1-error). In the parenthesis, we show its jackknife estimate of MCE. The SDs, ESDs, and Covs are averaged over the 15 elements in $\pi$.
  • Time is the average seconds taken for sampling 1,000 MCMC draws in Matlab on a desktop with a 4.9GHz processor and 64GB RAM. The study is based on $R=200$ replications. 2,000 MCMC draws are obtained for each replication. The average of the inefficiency factors is around 9.5 with standard deviation 1.4.
@@
Figure SE.3: Histograms of estimated posterior probabilities of $H_{1}$ in each of the 200 simulations under (a) case 1 (dependent consideration - $H_{1}$ is true) and (b) case 2 (independent consideration - $H_{0}$ is true). $\varepsilon=0.1,n=50$. $T=5$. With random effects.